• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name A. Nathan

    Log entry time 00:14:41 on January27,2002

    Entry number 80954

    keyword=Shift Report

    1. As the shift starts, we are in production running on kinematics 4H(e), run 1714.
    2. During period 1600-1700, 4H(e) runs 1714-1718 completed successfully. Starting run 1719.
    3. During period 1700-1800, 4H(e) runs 1719-1722 completed successfully.
    Run 1723 in progress.
    4. During period 1800-1900, we had DAQ problems (18:10-18:37). See halog 80914. Runs 1725-6 are bad. Good runs are 1723-1724 and 1727-1728, all 4H(e) production.
    5. During period 1900-2000, 4H(e) runs 1729-1733 completed successfully. Reboot ROC 14 and TS2 at 2000.
    6. During period 2000-2100, 4H(e) completed (1736-38) as well as optics run (1739)
    7. At 2110, we had a controlled access to move to 4G kinematics. The hall returned to power permit at 2145, a new record for the RCS group, and beam returned at 2200 (a total of 50 minutes from beam off to beam on).
    8. Optics run 1740 completed for 4G at 2212.
    9. LH2, radiator out, endpoint runs (mon and moff) completed by shift end
    ..very productive shift!